OpenStudy (starfireandrobin):

Can someone tell me how to divide 39.59 by 1.07? On the calculator, it says "37", but writing it on the notebook, the first number I get is "5." Can somebody explain?

OpenStudy (starfireandrobin):

@Vuriffy since I see you're online lol

OpenStudy (daniellelovee):

is 37

OpenStudy (starfireandrobin):

No yes, I know. But how do I do this on a piece of paper?

OpenStudy (starfireandrobin):

Can you write an equation for me? Or explain step by step?

OpenStudy (starfireandrobin):

I first made 1.07 into a whole number, moving the decimal point two spots to the right. Then I did the same to 39.59.

OpenStudy (tkhunny):

Don't do that. When working with two decimal places, why would you round to 0 decimal places before you even start? 39.59 / 1.07 = 37.00 It probably would help if you had a fixed number of decimal places in your display, rather than letting it pithc your significant digits.

OpenStudy (mhchen):



OpenStudy (tkhunny):

3959 / 107 = 37

OpenStudy (daniellelovee):

Shift the decimal points of both 39.59 and 1.07 to the right by 2 decimal places. (This is needed to convert 1.07 to 107, since long division must divide by a whole number.) \frac{3959}{107} ​107 ​ ​3959 ​​ 2 Set up the long division. 1 0 7 | 3 9 5 9 3 Calculate 395 ÷ 107, which is 3 with a remainder of 74. 3 1 0 7 | 3 9 5 9 3 2 1 7 4 4 Bring down 9, so that 749 is large enough to be divided by 107. 3 1 0 7 | 3 9 5 9 3 2 1 7 4 9 5 Calculate 749 ÷ 107, which is 7. 3 7 1 0 7 | 3 9 5 9 3 2 1 7 4 9 7 4 9

OpenStudy (daniellelovee):

hope that helps sorry for the long reply